OpenStudy (anonymous):

-4n+7+2n=1 n = 2.666666667 (this is what i got and i dont think it is right, none of my choices are decimals

zepdrix (zepdrix):

\(\large\rm -4n+7+2n=1\) Subtracting 7 from both sides gives us \(\large\rm -4n+2n=1-7\) Then combine like terms, \(\large\rm -2n=-6\) Any confusion on those two steps?

OpenStudy (anonymous):

so the answer would be -6 ??

zepdrix (zepdrix):

No, at this point we've solved for -2n, not n.

zepdrix (zepdrix):

To undo the 7 being `added` to our n, I `subtracted` 7 because that's the opposite of addition. n is being `multiplied` by -2. How do you undo multiplication?

zepdrix (zepdrix):

What's the opposite of multiplication?

OpenStudy (anonymous):

division

OpenStudy (anonymous):

wait, what am i gonna divide by ?

zepdrix (zepdrix):

Good. Since n is being multiplied by -2, we can undo that operation by division by -2.

zepdrix (zepdrix):

\[\large\rm \frac{-2n}{-2}=\frac{-6}{-2}\]On the left side of the equation, the multiplication of -2, and division of -2 undo one another,\[\large\rm n=\frac{-6}{-2}\]

OpenStudy (anonymous):

i am really confused.... how do i divide n ?

zepdrix (zepdrix):

You're dividing -2n by -2. Just divide the number in front of n by the number in the bottom.\[\large\rm \frac{-2n}{-2}\quad=\frac{n}{1}\]-2 divided by -2 is 1. So that leaves us with n/1 or simply n.

OpenStudy (anonymous):

oh oh oh i should have knew that